OpenStudy (anonymous):

1/2 divided by 1/4 divided by 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.5 (1/2) divided by .25 (1/4) divided by 8 = .25 ---> 1/4

OpenStudy (anonymous):

^^ I just converted the fractions into decimals. But there are other methods.

OpenStudy (anonymous):

so all i had to do was convert the fractions to decimals first? thanks so much!

OpenStudy (anonymous):

Sorry. Thanks. Hero's right. I meant .5 = 1/2 and .25 = 1/4

OpenStudy (anonymous):

its ok, i knew what you were saying